Consider the following probability distribution of the random variable X.

                                                               

x

2

4

6

8

10

12

P(Y = x)

0.05

0.15

0.25

A

0.05

0.25

State the value of A to two digits to the right of the decimal point.
Answer #1

Solution :

  P(y=x)=1

=0.05+0.15+0.25+A+0.05+0.25=1

=0.75+A=1

A=1-0.75=0.25

A=0.25